Water ionizes by the equation H2O(l)⇌H+(aq)+OH−(aq) The extent of the reaction is small in pure water and dilute aqueous solutions. This reaction creates the following relationship between [H+] and [OH−]: Kw=[H+][OH−] Keep in mind that, like all equilibrium constants, the value of Kw changes with temperature. Part A What is the H+ concentration for an aqueous solution with pOH = 3.02 at 25 ∘C?
Answer #1

Answer

[H+] = 1.05×10-11M

Explanation

[H+] [OH-] = Kw

Kw is called ionic product of water

Value of Kw at 25℃ is 1.00×10-14

pOH = - log[OH-]

-log[OH-] = 3.02

[OH-] = 9.55×10-4M

9.55×10-4M × [H+] = 1.00×10-14M2

[H+] = 1.05 × 10-11